OpenStudy (anonymous):

Multiply and simplify: (3x – 7)(3x + 7) A. 9x2 – 42x – 49 B. 9x2 – 42x + 49 C. 9x2 + 49 D. 9x2 – 49

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3x(3x + 7) - 7(3x + 7)}\)

OpenStudy (anonymous):

Use: \[(a-b)(a+b) = a^2 - b^2\]

OpenStudy (anonymous):

^ which is the Difference of Perfect Squares rule.

Parth (parthkohli):

Well you can use either methods

OpenStudy (anonymous):

Using Difference of Perfect Square you get: 9x^2-49 Therefore, answer is D.

OpenStudy (anonymous):

i have a simple formula\[(a+b)\times(a-b) =a ^{2} -b ^{2}\]

OpenStudy (anonymous):

so we get answer \[9x ^{2}-49\]